Systems Staff Meeting Notes 8/2/99

Systems Staff Meeting Notes
August 2, 1999

Present: Rich Bennett, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Announcements
  1. Rich had an excellent vacation - spent much of it going horseback riding, waterfall hunting and at an amusement park.
  2. Debra's band was featured in last week's Scene magazine.
  3. The Network Coordinator's position will have to be re-posted - our last candidate, Adam, had another job offer come his way.
  4. Brian Burton's old position has been posted. The pool should be available August 8.
  5. Bill Covey will be out Wednesday-Friday this week on an FCLA field trip to Atlanta to visit the Systems departments of Georgia Tech & Emory.
  6. Martha Hruska will be out of the office all this week.
  7. Ying will be moving into the office formerly known as Brian's, room 18C.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. A nasty security hole has been found in Office 97, which can permit an intruder to do pretty much anything to your machine. Unlike most previous holes, this one does not require you to open a hostile document to be activated; merely visiting the wrong Web page is sufficient. It also can be activated via email attachments and Office files. There is a security patch for Office'97 that needs to be installed immediately. We will put it out on the server for the liaisons to access. Office 2000 has the fix, but since it will be at least three weeks until we receive it, we have to install the security patch on all machines with Office'97. The download is fairly large.
  2. There are 10 Cows left to be picked up.
  3. A special order for Accounting will be going out soon. They'll be getting five new work stations. Two with big monitors - 20" or more.
Liaisons
  1. Liaison training program to start next week.
Networking
  1. Switches have been installed in West, hubs have been removed. There have been some management problems which will be fixed in the next software release in October '99.
  2. NERVM mail - POP server is gone - remains of mail server will go in October. IMAP still works here.
  3. The Proxy Server went down last Thursday, it ran out of space. Had a false report of it malfunctioning today.
Software Projects Status
  1. A change to Brian's LASSIE database has been requested - it will be put on hold until the new programmer is hired.
  2. Ying is in the process of making the appropriate changes to the IR databases following Bobbie's evaluation of the system.
  3. Alena has an editing option on her database now. It's not working at 100% - it starts very slowly then picks up after 2-3 records. Bill suggests that a new OBDC driver may help.
  4. Trouble Ticket Viewer problem has been eliminated.
  5. Mellon #1 - hopefully the last changes have been made on the database images.
DP Coordinator
  1. This week marks the end of the summer term. There will be no more students for the next 2 weeks.
  2. Debra is working on the new Fall schedule.
  3. Lindsay will be moving off the front desk and into Scott's position in the Order Room. Scott will not be working at all, he'll be concentrating solely on school.
Calendar
  1. Bill Covey will be out August 4-6.
  2. Martha Hruska is out August 2-6.
  3. Debra will be out August 16-20.


Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu